Write the equation of a line perpendicular to the given line but passing through the given point.
y= 2x -1
pass through points are 2 and -2
Also simplify equations into slope intercept form

Answer by mananth(16946)

y=2x-1
compare with y=mx+b
m=2 slope
perpendicular line will havea slope of -1/2
and passes through (2,-2)
use the slope point formula
y-y1= m(x-x1)
y-(-2)=-(1/2)(x-2)
y+2=-1/2x +1
y=(-1/2)x -1 is the equation